OpenStudy (anonymous):

simplify X^3/4 * 2/X^2

OpenStudy (mertsj):

\[x ^{\frac{3}{4}\times}\frac{2}{x^2}=\frac{2x ^{\frac{3}{4}}}{x^2}=2x ^{\frac{3}{4}-2}\]

OpenStudy (mertsj):

You can finish it.

OpenStudy (anonymous):

lets walk through it

OpenStudy (mertsj):

Which part?

OpenStudy (anonymous):

the end

OpenStudy (mertsj):

So you're having trouble with the subtraction of 3/4-2?

OpenStudy (anonymous):

no i got -1.25 like is that it?

OpenStudy (mertsj):

yes. That is the exponent